# Exploiting the Uncoordinated Privacy Protections of Eye Tracking and VR Motion Data for Unauthorized User Identification

### v1

@@ Abstract
Virtual reality (VR) devices use a variety of sensors to capture a rich body of user-generated data, which can be misused by malicious parties to covertly infer information about the user. Privacy-enhancing techniques seek to reduce the amount of personally identifying information in sensor data, but these techniques are typically developed for a subset of data streams that are available on the platform, without consideration for the auxiliary information that may be readily available from other sensors. In this paper, we evaluate whether body motion data can be used to circumvent the privacy protections applied to eye tracking data to enable user identification on a VR platform, and vice versa. We empirically show that eye tracking, headset tracking, and hand tracking data are not only informative for inferring user identity on their own, but contain complementary information that can increase the rate of successful user identification. Most importantly, we demonstrate that applying privacy protections to only a subset of the data available in VR can create an opportunity for an adversary to bypass those privacy protections by using other unprotected data streams that are available on the platform, performing a user identification attack as accurately as though a privacy mechanism was never applied. These results highlight a new privacy consideration at the intersection between eye tracking and VR, and emphasizes the need for privacy-enhancing techniques that address multiple technologies comprehensively.
@@ 1 Introduction
Virtual reality (VR) systems collect a large amount of data from a variety of sensors, including those that track the position and movement of a user’s body over time. The information captured by sensors such as Inertial Measurement Units (IMUs) track a user’s motion over time, which enables many of VR’s core functionalities. However, the collection and utilization of this data may also inadvertently reveal more information about a user than what is necessary for a given application. For example, while motion tracking data is useful for delivering a sense of immersion to the user, it also contains information about a user’s physical traits and behavioral patterns, which can be misused to infer more information about a user than they anticipated. The potential misuse of motion tracking data to reveal excessive amount of information about a user raises concerns for user privacy, as a user may be profiled in ways that they did not initially consent to or re-identified across play sessions without their knowledge.
@@
In addition to sensors that track the user’s body in a virtual space, eye tracking technology is becoming an increasingly standard feature on VR platforms. Eye tracking enables several vital utilities in VR, including the delivery of power savings through foveated rendering [26], facilitating the correction of the visual distortions that contribute to VR sickness [18], and enhancing the perceived realism of virtual avatars during social interactions [8]. Although it delivers these vital functionalities to VR platforms, eye tracking also complicates VR’s privacy landscape by introducing additional concerns for user privacy. Like motion data in VR, eye tracking data also captures sensitive information about a user that may be used to infer personal attributes [11]. Whereas motion data in VR may convey information about a user’s physical traits and kinesthetic behavior, eye tracking can reveal information about a user’s cognitive and behavioral traits, such as their personality or interests. If eye tracking data is misused to reveal sensitive characteristics about a user, these insights can complement the rich body of user information that can already be inferred from motion data.
While a growing body of research has been dedicated to investigating the privacy implications of eye tracking and VR motion data separately, relatively fewer works address the privacy concerns that may arise when these data streams are available simultaneously. Because motion tracking and eye tracking data can reveal different user characteristics, it is imperative to consider their combined potential to enable novel privacy violations in platforms where they are both available. To address this novel consideration, we explore how the simultaneous availability of these data streams can introduce new, unexplored privacy vulnerabilities when existing privacy-enhancing techniques fail to account for the combined risk of user identification across multiple sensors.
@@
In this work, we conduct an empirical study that builds on prior research demonstrating the potential for both VR motion data and eye tracking data to identify users on virtual platforms. Using a publicly available data set of eye tracking, VR headset, and VR controller motion data, we demonstrate that VR motion data and eye tracking signals can be used to identify users in VR platforms, and examine the extent to which the application of privacy-enhancing mechanisms can reduce the risk of user identification. We then simulate an adversarial privacy attack where data from VR motion sensors is used to circumvent the privacy protections applied to eye tracking data, and vice versa. Our results show that, without comprehensive privacy safeguards that address all available data sources, even state-of-the-art privacy mechanisms can be bypassed, allowing for high-accuracy user identification through unprotected data streams. This study contributes to a fundamental understanding of novel privacy considerations that exist at the intersection between VR and eye tracking, and provides motivation for future work developing comprehensive privacy solutions that are suitable for use in multi-sensor environments such as VR.
@@ 2 Background and Related Work
VR platforms capture user activity and environmental information using a variety of sensors, such as IMUs, proximity sensors, and cameras. While this data fulfills the benign purpose of enabling many of VR’s core functionalities, the user-generated portion of the data can also convey information about the user beyond what is strictly necessary for a VR application have knowledge of. For example, the average height of the headset in the world space can reveal the user’s approximate height [23]. The information that can be inferred about a user is not limited to physical characteristics about their body’s proportions; VR data can also be analyzed to infer secondary personal characteristics such as gender [30, 9, 22], age [30, 9, 22], and several other socioeconomic traits that are correlated with these characteristics [22]. Taken together, this information can be used to build a virtual profile of a user based on their motion data. Several works in literature demonstrate that users can be uniquely identified through motion data collected in VR—this motion data conveys information about the user’s physical characteristics [22] and behavioral patterns [32] that can be used to re-identify them across a variety of virtual environments [20]. Users can also be re-identified across several play sessions or uniquely identified among a population of other users [19]. The ability to covertly extract this kind of personally identifying information from VR data has significant implications for user privacy, even if it does not always attach a user’s real-world identity to their virtual persona.
Efforts to protect user privacy in VR often manifest as system-level security protections that prevent unauthorized access to user-generated data, such as encryption, firewalls, and intrusion detection [12, 5]. While these techniques can help prevent data from being accessed through unambiguously malicious means like exploiting security vulnerabilities, these techniques do not protect users from privacy violations that can be committed by applications using data that is willingly provided by users to potentially malicious entities. VR applications often require motion data from various body tracking sensors to enable basic functionality, but this data can be used to covertly extract or infer sensitive information about users [21]. Malicious actors may also collect personally sensitive data by observing or recording another user’s movement in a play space [19], which may not evoke enough suspicion to prompt action before the data has been collected. When data is streamed from a device’s sensors to a malicious application, it is possible for these privacy violations to take place before a user recognizes that their privacy may be at risk. Policy-based solutions that dictate how VR data should be used may also be insufficient for protecting users from such privacy violations, because VR applications may not adhere to the data management practices stated in their privacy policies. By analyzing the network traffic of 140 popular VR apps, Trimananda et al. [31] observe that 70 percent of data transmissions were inconsistent with their respective application’s privacy policy, and that the data that was transmitted contained personally sensitive information that could be used for user identification. This suggests that additional mechanisms are needed to protect user privacy when policy-based solutions fail.
@@
Developing privacy-enhancing techniques for user-generated data is one open research area that attempts to fulfill the need for data-based privacy enhancement. Rather than attempting to litigate access to (or utilization of) potentially sensitive data, these techniques apply data processing techniques to user-generated data streams to obfuscate personally identifying details before they are made accessible to third parties. Preliminary efforts by Sun et al. [29] suggest that differential privacy can be applied to obscure behavioral patterns that may identify a user; however, this investigation was only performed on simulated human activity. Wierzbowski et al. [32] observe that eliminating differences in users’ characteristics (e.g., by standardizing all users’ heights by seating them in a chair) while collecting VR motion data can reduce user identification rates; this places constraints on how users can interact with their devices, which makes this method difficult to apply in a practical setting. Nair et al. [23] apply differential privacy to obfuscate physical characteristics from existing data without restricting how users interact in VR. Similarly, Meng et al. [19] propose adding noise to user motion data before it is used to render the virtual avatar’s motion, thereby obscuring the user’s movement patterns.
@@
The privacy landscape for VR is further complicated when considering the growing integration of eye tracking technology in VR platforms. Eye tracking is used to provide vital functionalities to VR platforms, such as power savings via foveated rendering [26] and the correction of visual distortions that cause discomfort to users [18]. However, like other sensors that capture human activity, eye tracking also introduces concerns for user privacy by way of enabling user profiling and identification. Eye tracking data can reveal personal information about the user that may be used to profile and identify them. Eye movement alone can serve as a behavioral biometric that uniquely identifies users [17], and can be used to infer information about a user’s cognitive characteristics, such as their fatigue levels [34], personality [2], cognitive load [25], and emotional state [15]. Several characteristics that can also be inferred through VR, such as gender and age, can also be inferred through eye tracking data [11]. Efforts to enhance privacy in eye tracking include system-level techniques to restrict the transmission of sensitive data [10, 3, 4], and the application of data-augmenting privacy mechanisms to remove sensitive information from gaze data while allowing it to be used for benign purposes, for both real time gaze interaction [33, 13] and during post-hoc data analysis of previously collected gaze data [4, 28, 16].
@@
The privacy-enhancing techniques discussed thus far are proposed for either VR motion data or eye tracking data separately, under the assumption that no other potentially sensitive data streams are available to the adversary. While privacy protection for VR and eye tracking remain open areas of research on their own, it is becoming increasingly important to address the privacy implications that arise when they are used together. The joint integration of motion sensors and eye tracking on consumer-facing VR platforms (e.g, the Meta Quest Pro, Vive Pro Eye) is making the availability of both sources of data collected synchronously increasingly common, which can be problematic for user privacy if not acknowledged. These data streams can potentially provide complementary information that increase the accuracy of user identification; for example, an adversary may build a more complete profile of a user by combining complementary information about their physical bodies (extracted from VR motion) with information about the way they think, act, and allocate their attention (extracted from eye tracking data). Several works show that analyzing gaze behavior in conjunction with VR motion data can contribute to more successful user identification [14, 27, 32, 24]. Given that the privacy protection strategies for eye tracking and VR data have been developed in isolation from one another, it is possible that there are gaps in privacy research where the two technologies intersect. In this work, we explore one such potential gap where an adversary attempts to circumvent eye tracking privacy protections by using non-privatized VR data, and vice versa, to perform user identification.
@@ 3 Threat Model
For this investigation, we explore the problem of unauthorized user identification in VR platforms. The threat model for this work, illustrated in Figure 1, involves a relatively privileged adversary who attempts to perform user identification using the user-generated body motion and gaze data collected by various sensors on a VR device. This adversary does not necessarily need to engage in malicious activity to gain access to these streams; if they are a developer for an application in VR, they could be granted access to body telemetry and gaze data streams to enable the functionality of their otherwise benign application [21]. As such, security measures that detect unusual data transmission or unauthorized access to information would not necessarily prevent this attack from occurring.
Given access to this data, the adversary then attempts to infer user identities using the signals that they collected from the device. The adversary may or may not have knowledge of which data streams (if any) had privacy mechanisms applied to them before they were provided by the device. We explore whether this adversary would be able to successfully perform user identification when (a) none of the data streams have had privacy protections applied, (b) all of the data streams have had privacy protections applied, and (c) some, but not all, data streams had privacy protections applied.
@@ 4 Methodology
In this investigation, we are interested in exploring two dimensions of user identification in this problem space. We perform an empirical study where we attempt to perform user identification on a publicly available data set of eye movement data and body motion data collected in VR. Given that gaze data and body motion data is often collected synchronously for benign purposes, we then evaluate whether combining these data streams increases the accuracy of user identification. We then investigate whether these data streams could be used to circumvent privacy-preserving techniques meant to prevent user identification, when only a subset of data undergoes privacy enhancement.
@@ 4.1 Data Set
We utilized OpenNEEDS [6], a data set of gaze, head, and hand signals that were collected from 44 participants while they freely explored indoor and outdoor virtual environments in VR. This data was collected in two sessions, where participants were instructed to explore their virtual environment and interact with various objects at their discretion. During each session, gaze data was captured using a custom eye tracker embedded in a VR headset, and motion tracking data was captured for the head, left hand, and right hand using sensors and controllers for an Oculus Rift VR headset. All data was captured at rate of 90 Hz. We excluded participants from this investigation if they did not have at least two sessions’ worth of data, or if they had fewer than 15 seconds’ worth of data in either session. As a result, we utilized data from 38 participants total.
@@ 4.2 Privacy Mechanisms
We selected two privacy mechanisms that have been previously proposed and validated for this investigation; these mechanisms have demonstrated the ability to enhance user privacy for gaze and VR head motion data, respectively. However, the two have not been tested in conjunction with one another. These mechanisms were selected because they are feasible to implement in real-time and produced favorable privacy-utility trade-offs in their original publications, which are important requirements to ensure that privacy protections do not interfere with the system’s overall usability.
We are interested in investigating the potential for circumventing the strongest privacy settings explored in each paper, so we apply each privacy mechanism using strongest privacy setting that was recommended in each original work. While privacy settings are typically fine-tuned to suit the data set’s need for privacy and utility balance, we found the privacy budgets proposed by each method were appropriate for use with our data. In this section, we give a brief overview of each of these mechanisms.

### v3

@@ Abstract
Virtual reality (VR) sensors capture large amounts of user data, including body motion and eye tracking, that contain personally identifying information. While privacy-enhancing techniques can obfuscate this data, incomplete privacy protections risk privacy leakage, which may allow adversaries to leverage unprotected data to identify users without consent. This work examines the extent to which unprotected body motion data can undermine privacy protections for eye tracking data, and vice versa, to enable user identification in VR. These findings highlight a privacy consideration at the intersection of eye tracking and VR, and emphasize the need for privacy protections that address these technologies comprehensively.
@@ 1 Introduction
Virtual reality (VR) systems collect a large amount of data from a variety of sensors, including those that track the position and movement of a user’s body. Sensors such as Inertial Measurement Units (IMUs) track a user’s motion over time, which enables many of VR’s core functionalities. However, the collection and utilization of such data may inadvertently reveal more information about a user than what is necessary for a given application. For example, while motion tracking data is useful for delivering an immersive user experience, it also can also convey physical traits and behavioral patterns that may be used to infer additional information about the user. The potential misuse of motion tracking data to reveal an excessive amount of information about a user raises concerns for user privacy, as a user may be profiled or even re-identified across play sessions without their knowledge.
@@
In addition to sensors that track the user’s body in a virtual space, eye tracking technology is becoming an increasingly common feature on VR platforms. Eye tracking enables several functions in VR, including the delivery of power savings through foveated rendering [34], facilitating the correction of visual distortions that contribute to VR sickness [24], and improving the perceived realism of virtual avatars during social interactions [9]. In addition to delivering these functionalities to VR platforms, eye tracking introduces additional concerns for user privacy in VR platforms. Like motion data in VR, eye tracking data also captures sensitive information about a user that can be used to infer personal attributes [14]. However, whereas motion data in VR may convey information about a user’s physical traits and kinesthetic behavior, eye tracking can more directly reveal a user’s cognitive behavioral traits such as their personality or interests. If eye tracking data is misused to reveal sensitive characteristics about a user, these insights can complement the rich body of user information that can already be inferred from VR motion data.
While much research has been dedicated to investigating the privacy implications of eye tracking and VR motion data separately, relatively fewer works address the privacy implications that may arise when these data streams are available simultaneously. The combination of these data types could increase the potential for privacy violations in ways that have not yet been fully understood.
@@
To address this gap, we explore how the simultaneous availability of these data streams can introduce privacy vulnerabilities, especially when existing privacy-enhancing techniques fail to account for the combined risk of user identification across multiple sensors. Using a publicly available data set of eye tracking, VR headset, and VR controller motion data, we demonstrate that VR motion data and eye tracking signals can be used to identify users in VR platforms, and examine the extent to which the application of privacy-enhancing mechanisms to this data can reduce the risk of user identification. We then simulate an adversarial privacy attack where data from VR motion sensors is used to circumvent the privacy protections applied to eye tracking data, and vice versa. Our results show that without comprehensive privacy safeguards addressing all available data sources, these privacy mechanisms can be bypassed to enable user identification through unprotected data streams. These findings underscore the need for comprehensive privacy solutions that can account for the novel privacy risks of multi-sensor environments in VR. This study contributes to a deeper understanding of how the privacy considerations for VR motion data and eye tracking data interact, and provides motivation for future work that develops comprehensive privacy solutions that are suitable for use in multi-sensor VR environments.
@@ 2 Background and Related Work
When using a VR platform, user activity is captured and recorded by a variety of sensors. VR platforms capture user activity using a variety of sensors, including IMUs, proximity sensors, and cameras. In addition to fulfilling benign functionalities, this activity data can also convey information about the user that is not strictly necessary for a VR application to have knowledge of; for example, the average height of the headset in the user’s world space over time can reveal the user’s approximate height [31]. The information that can be inferred about a user is not limited to physical characteristics like body proportions; VR data can also be analyzed to secondarily reveal personal characteristics such as gender [38, 11, 30] age [38, 11, 30], and socioeconomic traits that are correlated with these characteristics [30]. Taken together, this information can be used to build a virtual profile of a user based on their motion data.
Several works in literature demonstrate that users can be uniquely identified through VR motion data—this data conveys information about the user’s physical characteristics [30] and behavioral patterns [40] that can be used to re-identify them across play sessions in a virtual environment [26]. Users can also be re-identified across several play sessions or uniquely identified among a population of other users [25]. The ability to covertly extract personally identifying information from VR motion data has significant implications for user privacy, even if a user’s real-world identity is not always attached to their virtual persona. Although a privacy vulnerability of this kind is concerning in any application, it becomes particularly problematic in emerging VR use cases where data is gathered from sensitive populations, such as children [4], or applied in privacy-sensitive contexts such as delivering healthcare [16] or education [13].
Efforts to protect user privacy in VR often manifest as system-level security protections that prevent unauthorized access to user-generated data, such as encryption, firewalls, and intrusion detection [15, 6]. Although these techniques can help prevent data from being accessed through unambiguously malicious activities such as exploiting security vulnerabilities, these techniques do not prevent privacy violations that can be committed using data that is willingly provided by users. When VR motion data is provided to a malicious application, it is possible for these privacy violations to take place before the privacy risk is recognized by either the system or the user. For example, a VR application that is granted motion data for a benign purpose may secondarily use it to infer sensitive information about the user. Nair et al. [29] explore one such scenario by designing a VR game that obstinately collects motion data to enable gameplay, but which covertly elicits motion data that reveals sensitive user information.
Policy-based solutions that dictate how VR data should be used may also be insufficient for protecting users, because VR applications may not adhere to the data management practices stated in their privacy policies. After analyzing the network traffic of 140 popular VR applications, Trimananda et al. [39] observe that 70 percent of data transmissions were inconsistent with their respective application’s privacy policy, and often transmitted personally sensitive information that could be used for user identification. These gaps in security-based and policy-based solutions suggest that additional mechanisms are needed to protect users from being profiled from their VR motion data.
@@
To fulfill the need for privacy enhancement in VR, several works have proposed privacy-enhancing techniques that make personally identifying information more difficult to extract from user-generated data. Rather than litigating access to (or utilization of) potentially sensitive data, these techniques apply data processing techniques to user-generated data streams to obfuscate personally identifying details before they are made accessible to third parties. Preliminary efforts by Sun et al. [37] suggest that differential privacy can be applied to obscure behavioral patterns that may identify a user; however, this investigation was only performed on simulated human activity. Wierzbowski et al. [40] observe that eliminating differences in users’ characteristics (e.g., by standardizing all users’ heights by seating them in a chair) while collecting VR motion data can reduce user identification rates; this places constraints on how users can interact with their devices, which makes this method difficult to apply in a practical setting. Nair et al. [31] apply differential privacy to obfuscate physical characteristics from existing data without restricting how users interact in VR. Similarly, Meng et al. [25] propose adding noise to user motion data before it is used to render the virtual avatar’s motion, thereby obscuring the user’s movement patterns in a virtual space.
@@
The privacy landscape for VR is further complicated when considering the growing integration of eye tracking technology in VR platforms. Eye tracking is used to provide vital functionalities to VR platforms, such as power savings via foveated rendering [34] and correction of visual distortions that can cause discomfort to users [24]. However, like other sensors that capture human activity, eye tracking also introduces concerns for user privacy by enabling user profiling and identification. Eye tracking data is a particularly rich source of information that gives insights into a user’s cognition. It can serve as a behavioral biometric that uniquely identifies users [22], and can be used to infer information related to the user’s cognitive state, such as their fatigue level [42], cognitive load [33], and emotional state [20] and personality type [2]. Several personal characteristics that can be inferred through VR, such as gender and age, can also be inferred through eye tracking data [14].
As eye tracking becomes an increasingly standard feature in consumer-facing VR devices—including Meta’s Quest Pro, HTC’s Vive Pro Eye, and Apple’s Vision Pro—privacy concerns regarding the use of this data have prompted the development of privacy-enhancing techniques for gaze data on these devices. Efforts to enhance privacy in eye tracking include system-level techniques to restrict the transmission of sensitive data [12, 3, 5], and the application of data-augmenting privacy mechanisms to remove sensitive information from gaze data while allowing it to be used for benign purposes, for both real time gaze interaction [41, 17] and during post-hoc data analysis of previously collected gaze data [5, 36, 21].
@@
The privacy-enhancing techniques discussed thus far were developed for VR motion data or eye tracking data in isolation, under the assumption that no other potentially sensitive data streams are available to the adversary. Although the privacy implications of VR and eye tracking remain open areas of research on their own, it is becoming increasingly important to address the privacy implications that arise when they are used together. The joint integration of motion sensors and eye tracking on consumer-facing VR platforms is making the availability of both sources of data collected synchronously increasingly common, which can be problematic for user privacy if not acknowledged. These data streams can potentially provide complementary information that increase the accuracy of user identification; for example, an adversary may build a more complete profile of a user by combining complementary information about their physical bodies (extracted from VR motion) with information about the way they think, act, and allocate their attention (extracted from eye tracking data). Several works show that analyzing gaze behavior in conjunction with VR motion data can contribute to more successful user identification [19, 35, 40, 32], but this trend has not been discussed in the context of its implications for user privacy.
Given that privacy protection strategies for eye tracking and VR data have been developed in isolation, additional privacy considerations may emerge when these technologies intersect. In this work, we explore one such consideration where an adversary attempts to identify users on a VR platform, circumventing incomplete privacy protections by leveraging data streams collected by sensors for VR motion tracking and eye tracking. While the potential privacy considerations of combining sensor data in VR platforms has been discussed conceptually in prior research [28, 10], this investigation empirically explores the impact of applying privacy mechanisms on user identification rates when VR motion and eye tracking data are available simultaneously. First, we empirically evaluate the effectiveness of two state-of-the-art privacy mechanisms designed for eye tracking and VR motion data in mitigating privacy leakage from this data, measured by the reduction of user identification rates. We also explore how incomplete privacy protections affect user identification rates by simulating a plausible scenario where privacy protections are applied only to a subset of the available data streams; in doing so, we demonstrate that incomplete privacy protection can enable unauthorized user identification.
@@ 3 Threat Model
For this investigation, we explore the problem of unauthorized user re-identification in VR platforms. The threat model for this work, illustrated in 1, involves an adversary who attempts to perform user identification using body motion and gaze data collected by a VR device. This adversary does not necessarily engage in malicious activity to gain access to these data streams; for example, if they are a developer for an application in VR, and could be granted access to body telemetry and gaze data streams to enable the functionality of their otherwise benign application [29]. As such, security measures that detect unusual data transmission or unauthorized access to information would not necessarily prevent this attack from occurring.
Given access to this data, the adversary then attempts to infer user identities using the signals that they collected from the device. The adversary may or may not have knowledge of which data streams (if any) had privacy mechanisms applied to them before they were provided by the device. Through repeated experimentation, they may even be able to identify which data streams are unprotected, and use this information to inform future attacks. We explore whether this adversary would be able to successfully perform user identification when (a) none of the data streams have had privacy protections applied, (b) all of the data streams have had privacy protections applied, and (c) when some, but not all, data streams had privacy protections applied.
@@ 4 Methodology
Given that gaze data and body motion data is often collected synchronously for benign purposes in VR, we first seek to evaluate whether combining these data streams increases the accuracy of user identification. We then investigate whether these data streams could be used to circumvent privacy-preserving techniques meant to prevent user identification, when only a subset of data undergoes privacy enhancement. We performed an empirical study where we attempted user identification on a publicly available data set of eye movement and body motion data collected in VR.
@@ 4.1 Data Set
We utilized OpenNEEDS [7], a data set of gaze, head, and hand signals that were collected from 44 participants while they freely explored indoor and outdoor scenes in VR. This data was collected in two sessions, where participants were instructed to explore their virtual environment and interact with various objects at their discretion. During each session, gaze data was captured using a custom eye tracker embedded in a VR headset, and motion tracking data was captured for the head, left hand, and right hand using equipment from an Oculus Rift. All data was captured at a sampling rate of 90 Hz. We excluded participants from this investigation if they did not have at least two sessions’ worth of data, or if they had fewer than 15 seconds’ worth of data in either session. As a result, we utilized data from 38 participants total.
@@ 4.2 Privacy Mechanisms
We selected two privacy mechanisms that have been previously proposed in literature for this investigation; these mechanisms have demonstrated the ability to enhance user privacy for gaze and VR head motion data, respectively. However, the two have not been tested in conjunction with one another. These mechanisms were selected because they are feasible to implement in real-time and produced favorable privacy-utility trade-offs in their original publications, which are important requirements to ensure that privacy protections do not interfere with the system’s overall usability.
We applied each privacy mechanism using the strongest privacy settings recommended in each paper. While privacy settings are typically fine-tuned to suit an individual data set, we found that the privacy budgets proposed by each method were appropriate to use with our data. In this section, we give a brief overview of each of these mechanisms.